OpenStudy (dj3strella):

3. What value should be added to the expression to create a perfect square? x2 – 10x 5 25 100 –100

OpenStudy (jdoe0001):

have you covered "perfect square trinomials" yet?

OpenStudy (dj3strella):

Yes @jdoe0001

OpenStudy (anonymous):

Rule of thumb of perfect squares is c=(b/2)^2 for the form ax^2+bx+c

OpenStudy (anonymous):

ex. x^2+20x+? c=(20/2)^2=10^2=100 c=100

OpenStudy (jdoe0001):

\(\bf x^2 - 10x\implies x^2 - (2\cdot 5)x\implies x^2 - 2( 5)x \\ \quad \\ thus\quad x^2 - 2( 5)x+{\color{red}{ \square }}^2\) any ideas?